OpenStudy (anonymous):

PLEASE HELP D:

OpenStudy (anonymous):

how do I write this equation: V= (12)(8)(2) = 192 inches^3 in terms of x?

OpenStudy (anonymous):

Please walk me through step by step don't just give me the answer....

OpenStudy (mertsj):

How can you write it in terms of x if it does not contain an x?

OpenStudy (anonymous):

I'm not sure :/

OpenStudy (anonymous):

The directions phrase it like this exactly.... "The equation for the volume of the object written in terms of x...."

OpenStudy (mertsj):

Could you post the entire question?

OpenStudy (anonymous):

is there an image with this problem?

OpenStudy (anonymous):

no there isn't a image

OpenStudy (anonymous):

That was the entire question. This is part of a project I'm doing for Algebra two and I have to solve my partners demsions

OpenStudy (mertsj):

So I take it this was a lab?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

Could you post the directions?

OpenStudy (anonymous):

Directions: A description of the rectangular box chosen My partner chose a rectangular box. The length, width, height and volume of the object The dimensions of my partner’s rectangular object is length 12 inches, width 8 inches and height 2 inches with a volume of 192 inches^3. The equation for the volume of the object written in terms of x The graph of the function The solutions of the equation including the algebraic work used to find the solutions.

OpenStudy (phi):

The instructions should mention what they mean by "x". Read over the problem carefully and look for any mention of "x".

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It just says to write it in this form.... x(x ) (x )

OpenStudy (phi):

It sounds like they want you to write the length, width and height as x, (x+a) and (x+b) (for example). Do they say x is the width, length or height? Or can you choose any of the 3 to be x ?

OpenStudy (anonymous):

No it says x is the (wlh) not to choose any of the 3

OpenStudy (phi):

Take a screen shot of the instructions and post them here.

OpenStudy (anonymous):

won't let me paste it here :/

OpenStudy (anonymous):

Hold on!

OpenStudy (phi):

What does the first line of 3) say ?

OpenStudy (anonymous):

It says: "Rewrite the formula using the variable x for the length. Substitute the value of the volume found in step 2 for V and express the width and height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the length, then the expression for the height will be (x + 4). "

OpenStudy (phi):

Just the first sentence. what does it mean ***using the variable x for the length***

OpenStudy (anonymous):

Uh...idk

OpenStudy (phi):

It means "let x be another name for the length" what is the length for your box ?

OpenStudy (anonymous):

well its not the length of my boy im trying to find it's my partner's box (cause we had to find each other's box)

OpenStudy (anonymous):

The length of his is 12 inches

OpenStudy (phi):

so call the length x, x = 12 what is the width ?

OpenStudy (anonymous):

Width = 8in

OpenStudy (phi):

now the interesting part: we want to write width= (x+a) width is 8, x is 12, so you know enough to figure out what a is. 8 = 12 + a what is a ?

OpenStudy (anonymous):

4

OpenStudy (phi):

are you saying 8 = 12+4 ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

is that what you calculator says?

OpenStudy (anonymous):

no I didn't use a calculator

OpenStudy (anonymous):

I'm supposed to use a calculator for this? O.o

OpenStudy (phi):

No, you should be able to do this without a calculator. But 12+4 is not 8

OpenStudy (anonymous):

I'm confused :s

OpenStudy (phi):

you want a number that when we add it to 12 we get 8 the number has to be a negative number

OpenStudy (anonymous):

oh so negative 4?

OpenStudy (phi):

yes, 12-4= 8 (if you can't do this, you can at least use a calculator to check... but I would learn how to do it with out a calculator) so far we have length= x width= (x-4) what will the height be ?

OpenStudy (anonymous):

h = 2in

OpenStudy (phi):

yes, h= 2, but you want to write it like this: (x+a)= 2 x is 12, what must a be to get 2?

OpenStudy (anonymous):

your saying 2-a = ?

OpenStudy (phi):

no. you want height = (x+a) you need to find the negative number "a" that makes (x+a)= h

OpenStudy (anonymous):

so what + a = 2?

OpenStudy (phi):

x, (and remember x is the length, and the length is 12)

OpenStudy (anonymous):

oh ok so 12 + a = 2 ok so a = -10

OpenStudy (phi):

yes (x-10) = 2 so you can replace height (which is 2) with (x-10) your volume is 192= L*W*H but you can replace L with x, W with (x-4) and H with (x-10) your new equation is 192= x (x-4) (x-10)

OpenStudy (anonymous):

OMG THANK YOU FOR YOUR HELP!!

OpenStudy (phi):

that is the answer for part 3. For part 4, you must multiply out (use FOIL if you know it). You should post part 4 as a new question, so people can help you if you get stuck. Good luck, but I have to log off for a while...

OpenStudy (anonymous):

That's exactly what i was about to do and np! Thanks again! :D